# Amnesiac Flooding: Easy to break, hard to escape

### v1

@@ 3. Model and Notation
Unless stated otherwise we do not assume that nodes have access to unique identifiers, however they have locally labelled ports that are distinguishable and totally orderable for both receiving and sending messages. When we do work with identifiers these identifiers are assumed to be unique and drawn from $[|V|+\kappa]$ for $\kappa>0$ a constant. However, we assume that individual nodes have access to arbitrarily powerful computation on information they do have.
@@
Further, for $H$ a subgraph of $G$, we denote by $S_{H}$, $S$ restricted to $H$. We will refer to $u$ as the head of the message and $v$ as its tail. Below, we define the operator $A_{I,G}:2^{V^{2}}\to 2^{V^{2}}$ to implement one round of Amnesiac Flooding on the given configuration:
@@ 4. Uniqueness
Given the known sensitivity of Amnesiac Flooding to even minor faults and instances of asynchrony, a natural question is whether there exists a more robust protocol which preserves its desirable properties. Specifically, does there exist a terminating protocol for broadcast which obeys all of the following for all graphs and port labellings:
Conditions 4.1.
@@
The answer is strongly negative, in fact there is no other terminating broadcast protocol preserving these properties at all. We will actually prove the slightly stronger case, that this holds even if agents are provided with unique identifiers and are aware of the identifiers of their neighbours. Any broadcast protocol meeting the constraints of conditions 4.1 must be expressible in the following form:
@@
For a protocol $P=(b,f)$, a set $S\subset\mathbb{N}$ and a degree $k\in\mathbb{N}$ we describe $P$ as AF up to degree $k$ on $S$ if there is no graph of maximum degree $k$ or less labelled with only members of $S$ which distinguishes $P$ from $P_{AF}$. From here on we will assume that all unique labels are drawn from $[n+\kappa]$ where $\kappa$ is a sufficiently large constant and $n$ is the number of nodes in the graph. We obtain the following result, the proof of which makes up the remainder of this section, provided $\kappa\geq R(9,8)$.
@@
Let $P=(b,f)$ be a terminating broadcast protocol that meets the conditions of observation 4.1, then $P$ is not distinct from Amnesiac Flooding.
@@
The basic argument is to show that any terminating broadcast protocol that meets the criteria of 4.1 is identical to Amnesiac Flooding. Our core technique is to construct a set of network topologies such that any policy distinct from Amnesiac Flooding fails on at least one of them.
Through the use of a combinatoric argument we establish that for $\kappa\geq R(9,8)$, there must exist a set $T$ of at least $9$ identifiers with the following property: for all $u,v\in T$ if $u$ has only the neighbour $v$, upon receiving a message $u$ will not return the message to $v$. From here, we show that there must exist $U\subseteq T$ containing at least $6$ identifiers such that $P$ is AF up to degree $2$ on $U$. By constructing a set of small sub-cubic graphs, we are able to extend this to degree $3$.
This forms the base case of a pair of inductive arguments. First, we construct a progression of sub-cubic graphs which enforce that if $P$ is AF on $[m]$ up to degree $3$ it must be AF on $[m+1]$ up to degree $3$. We then construct a family of graphs which have a single node of high-degree and all other nodes have a maximum degree of $3$. Therefore, by the previous construction, all other nodes must behave as though running Amnesiac Flooding. These graphs permit an inductive argument showing that this unique high degree node must also behave as through running Amnesiac Flooding. In combination, these two constructions enforce that $P$ behaves like Amnesiac Flooding in all possible cases. The full proof is given in appendix B. ∎
5. Uniqueness: Relaxing the constraints
Despite the uniqueness established in the previous section. By even slightly relaxing any of the $4$ properties presented in Conditions 4.1, we are able to obtain correct and terminating protocols distinct from Amnesiac Flooding. We summarise this in the following result:
@@ 5. Uniqueness: Relaxing the constraints
For each of the constraints in conditions 4.1, there exists a relaxation permitting a correct and terminating protocol distinct from Amnesiac Flooding.
While the majority of these protocols are not practical, and most are similarly fault sensitive to Amnesiac Flooding, they demonstrate that a broad spectrum of protocols exist. Specifically, the relaxations we make use of are as follows:
  • Strict Statelessness: Several relaxations of this already exist, such as Stateless Flooding (initiator retains information for one round) or even classical non-Amnesiac Flooding (nodes are able to retain 1-bit for one round). We present a third relaxation, where nodes know the ID of their neighbours and their neighbours’ neighbours.
  • Obliviousness: We permit one bit of read only header information to be included with the message and used in routing decisions.
  • Determinism: We give each node access to one bit of randomness per round.
  • Unit Bandwidth: We allow each agent to send up to two messages per edge per round.
Since each of these constitutes only a minor relaxation of the restrictions, we argue that the uniqueness of Amnesiac Flooding is in some sense tight. We will present protocols and demonstrate their correctness and termination for each of these cases independently in appendix C.
@@ 6. Fault Sensitivity
A faux-even cycle (FEC) is a graph comprised of either two node disjoint odd cycles connected by a path or two odd cycles sharing only a single node. We denote by $FEC_{x,y,z}$ the $FEC$ with one cycle of length $12x+$ 1, one of length $12z+$ 1 and a path containing $y$ edges between them. We emphasize that if $y=0$ the two cycles share a common node and if $y=1$ the two cycles are connected by a single edge.
@@
Let $F=(V,E)$ be $FEC_{x,y,z}$, the even cycle representation of $F$ denoted $F_{2}$ is the graph constructed by splitting the end points of the interconnecting path in two, and duplicating the path to produce an even cycle. Formally, if the two cycles are of the form $a_{0}...a_{2x}a_{0}$ and $c_{0}...c_{2z}c_{0}$, with $a_{0}$ and $c_{0}$ connected by the path $b_{1}...b_{y-1}$, we construct the following large even cycle from four paths
@@
We first establish that balance is conserved by Amnesiac flooding and, as the empty configuration is balanced, Amnesiac Flooding cannot terminate from any imbalanced configuration. For Amnesiac Flooding not to terminate it requires that some message travels around the communication graph and returns to the same edge, in the same direction. We show that if a configuration is balanced, the trajectory of any message can spend only a bounded number of consecutive steps on any given cycle or FEC. These results allow us to show that no message’s trajectory that includes the same edge in the same direction twice could have begun from a balanced configuration. Thus, Amnesiac Flooding started from any balanced configuration must terminate. The full proof is given in appendix A. ∎
6.1. Consequences of Theorem 6.1
@@
Note that we refer to the Byzantine failures as weak, since they are only transient and only interfere with the forwarding of the message, not its content. It is obvious to see that in a stateless setting there is no way to deal with a Byzantine fault that changes the message as there is no way to verify which message is authentic. Intuitively, in our setting, Weak-Byzantine agents may choose to send messages to an arbitrary set of neighbours in each round and they are all controlled by a single coordinated adversary. We say that a Weak-Byzantine adversary with control of a given set of nodes can force some behaviour if there exists any weak byzantine failure on that set of nodes producing the forced behaviour. We can now express our fault sensitivity results, the proofs of which we defer to appendix D and begin with an extreme case of single message dropping.
Theorem 0 (Single Message Failure).
If single-source Amnesiac Flooding experiences a single message drop failure for the message $(u,v)$ then it fails to terminate. if either:
  • $uv$ is not a bridge
  • $uv$ lies on a FEC subgraph
Moreover, it fails to broadcast if and only if this is the first message sent along $uv$, $uv$ is a bridge, and the side of the cut containing $u$ does not contain an odd cycle.
Thus, Amnesiac Flooding is extremely non-fault tolerant with respect to message dropping.
@@
Theorem 0 (Uni-directional link failure).
@@
Theorem 0 (Byzantine Failure).
@@
  • Non-termination if and only if at least one member of $J$ lies on either a cycle or FEC subgraph.
@@ 7. Conclusions and Future Work
We also obtain an understanding of the structural properties of Amnesiac Flooding. In particular, we study its sensitivity to single message drops, uni-directional link failures, and weak byzantine collusion, showing it can easily become non-terminating or non-broadcasting under such conditions. This is, perhaps, surprising as statelessness is frequently associated with fault tolerance, such as in the self stabilizing setting. A reasonable interpretation of Theorem 6.1, however, is that Amnesiac Flooding, while locally stateless, depends heavily on a distributed “meta-state” contained in the configuration of sent messages. This suggests it is unlikely that any minor modification of Amnesiac Flooding will resolve its fragility without depending on an entirely different mechanism for termination. In support of this, we note that of the four alternatives presented in the proof of the Existence of Relaxed Protocols Theorem, only Random-Flooding is meaningfully more robust (and will in fact terminate from any configuration in finite time almost surely). Nevertheless, we contend that further exploration of stateless algorithms such as Amnesiac Flooding, their properties and related models are important for both theory and practice of distributed networks.
Acknowledgements.
We would like to acknowledge Danial Maqbool, Durham University, for useful related work and discussion as a summer intern.
@@ Acknowledgements.
Appendix A Proof of theorem 6.1
In this appendix we prove theorem Theorem 6.1. We begin by establishing that (im)balance is preserved by the operation of Amnesiac Flooding,
@@
Let $G=(V,E)$ with $H\subseteq G$ an odd cycle of length $12k+$ 1. If $m$ is a message of $S$ on $H$ with a message path of length $22k+$ 2 restricted only to $H$, then $S$ is imbalanced on $G$.
@@
For the sake of contradiction assume that $S$ is balanced on $G$ and $m\in S$ has a message path from $S$ of length $22k+$ 2 restricted only to $H$. Without loss of generality, assume that $m$ is travelling clockwise. On a cycle a message can collide with exactly one other message, upon doing which all of its message paths must terminate. If there is a message $m^{\prime}$ on $H$ travelling anti-clockwise in $S$, then either it collides with $m$ within $2k$ steps or it collides with another message travelling clockwise first. Since $S$ is balanced on $G$ there must be an equal number of messages travelling clockwise and anti-clockwise on $H$, and a collision removes one of each. Therefore, unless new messages are added to $H$ in subsequent steps, $m$ must collide with a message $m^{\prime}$ travelling anti-clockwise before it takes $12k+$ 1 steps. Thus, new messages must be added to $H$ that collide with the message $m$ would otherwise collide with. The introduction of a new message to $H$ must produce one message travelling clockwise and one travelling anti-clockwise, $m_{c}$ and $m_{a}$ respectively. Thus, $m_{c}$ must collide with $m^{\prime}$ before $m$ does and so must be introduced between $m$ and $m^{\prime}$. However, this means that $m$ will collide with $m_{a}$ before it would have reached $m^{\prime}$ giving our contradiction. ∎
@@
    $W$ cannot return to the node it just came from, i.e. no sequence of the form $uvu$.
@@
    $W$ cannot take $22x+$ 2 consecutive steps around an odd cycle of length $12x+$ 1 (Lemma A.5).
@@
If $W$ contains a factor $u...u$ then that factor contains an odd cycle as a subfactor.
@@
Assume there are at least two for contradiction. We take the first two cycles of odd length in $W$, they are either sequential ($u...u...v...v$), interleaved ($u...v...u...v$) or repetitive ($u...u...u$). In the sequential case this is a fully traversed $FEC$ and so in violation of Condition $4$. In the interleaved case, we have three subsequences $w_{1},w_{2},w_{3}$ of length $a,b,c$ respectively such that $a+b$ is odd and $b+c$ is odd (see Figure 2). Thus, $a+c$ is even. Since $w_{1}$ and $w_{3}$ induce an even cycle, either $a=c$ or $W$ is in violation of Condition $3$. Therefore, we have that $W$ travels exactly half way around an even cycle along $w_{3}$. Thus there must exist a message path $W^{\prime}$ that contains the factor $w_{1}w_{2}w_{1}$ which is more than $a+b$ steps around an odd cycle of length $a+b$ and so is in violation of Condition $2$. In the repetitive case, this is multiple traversals of an odd cycle in full and so also in violation of Condition $2$. Thus, there exists at most one odd cycle in $W$. ∎
@@
Claim: $C$ is an odd cycle.
If $C$ is of even length (say $2k$), then there must be an excursion from $C$ otherwise Condition $3$ would be violated. However, there can be at most one such excursion as it must contain a consecutive odd cycle and so the two subsequences on either side must be factors of $W$. Therefore, since some subsequence of $W$ traverses $C$ fully with one additional step, one of the two factors must take $k+1$ steps around $C$. This violates Condition $3$ and so $C$ must be an odd cycle.
Claim: $W$ consists of two odd cycles $C$ and $\hat{C}$ connected by a path.
If $C$ is an odd cycle there must be an excursion from it or Condition $2$ would be violated. This excursion must contain exactly one consecutive odd cycle which we denote by $\hat{C}$. If $\hat{C}$ does not share its starting node with $C$, $W$ either forms a path between $C$ and $\hat{C}$ or some chain of cycles. We can use Condition $5$ to eliminate all even cycles of this chain, then if there are any odd cycles in the chain they correspond to a fully traversed $FEC$ when paired with $\hat{C}$ and so violate Condition $4$. Thus, we have a $W$ that takes a simple path from $C$ to $\hat{C}$ and back, although possibly intersecting $C$ along the way.
Claim: Neither $C$ nor $\hat{C}$ intersect with the path between them.
Assume that the path to $\hat{C}$ does intersect $C$, since we are taking the same path in both directions any node shared between the path and $C$ appear in $W$ three times. This violates Lemma A.12 and so $C$ must be disjoint from the path to $\hat{C}$. Similarly $\hat{C}$ must be disjoint from the path otherwise it would violate the same lemma.
Claim: $C$ and $\hat{C}$ intersect.
If $C$ is disjoint from $\hat{C}$ the pair would form a fully traversed $FEC$, thereby violating Condition $4$. Thus, $C$ and $\hat{C}$ must intersect.
@@
Let $W=uv..wx_{1},...,x_{k}w..uv$ where $C=u...w...uv$ and the excursion to $\hat{C}$ is given by $wx_{1}...x_{k}w$. Now assume that $\hat{C}$ contains a node from $C$ which occurs strictly before $w$ in $W$. This node is on a consecutive odd cycle and appears twice. There must exist a latest such node in the ordering of $C$, we denote it $y$. Since $y$ is on a consecutive odd cycle and appears twice in $W$ it must be the start and end point of $\hat{C}$. However, then either $y=w$ which is impossible or $y$ appears three times, violating Lemma A.12. The same argument holds taking the earliest node shared by $C$ and $\hat{C}$ strictly after $w$. Thus, the only node that can be shared by $C$ and $\hat{C}$ is $W$, implying that $W$ forms two odd cycles sharing a single node. However, this is an $FEC$ which is fully traversed and so violates Condition $4$.
Thus, $W$ cannot exist and so there can be no recurrent message in $S$. ∎ We can now immediately obtain a proof of Corollary 6.2.

### v2

@@ 3. Model and Notation
Unless stated otherwise we do not assume that nodes have access to unique identifiers, however they have locally labelled ports that are distinguishable and totally orderable for both receiving and sending messages. When we do work with identifiers these identifiers are assumed to be unique and drawn from $|V|+\kappa]$ for $\kappa>0$ a constant. However, we assume that individual nodes have access to arbitrarily powerful computation on information they do have.
@@
Further, for $H$ a subgraph of $G$, we denote by $_{H}, SS$ restricted to $H$. We will refer to $u$ as the head of the message and $v$ as its tail. Below, we define the operator $A_{I,G}:2^{V^{2}}\to 2^{V^{2}}$ to implement one round of Amnesiac Flooding on the given configuration:
@@ 4. Uniqueness
In this section, we investigate broadcast protocols similar to Amnesiac Flooding and establish four desirable properties that Amnesiac Flooding uniquely satisfies in combination. On the other hand, we show that this result is sharp and that by relaxing any of these conditions one can obtain similar terminating broadcast protocols.
4.1. Uniqueness
Our first major result concerns the uniqueness of Amnesiac Flooding. Given the algorithms surprising properties, a natural question is whether other broadcast algorithms exist maintaining these properties. Specifically, does there exist a terminating protocol for broadcast which obeys all of the following for all graphs and valid port labellings:
@@ 4.1. Uniqueness
The answer is strongly negative. In fact there is no other terminating broadcast protocol preserving these properties at all. We will actually prove the slightly stronger case, that this holds even if agents are provided with unique identifiers and are aware of the identifiers of their neighbours. Intuitively, the Strict Statelessness condition forces any broadcast protocol to make its forwarding decisions based only on the messages it receives in a given round. The combination of Obliviousness and Unit Bandwidth forces any protocol meeting the conditions to view messages as atomic. Finally, Determinism forces the protocol to make identical decisions every time it receives the same set of messages. Formally, any broadcast protocol meeting the four conditions must be expressible in the following form:
@@
For a protocol $P=(b,f)$, a set $S\subset\mathbb{N}$ and a degree $k\in\mathbb{N}$ we describe $P$ as AF up to degree $k$ on $S$ if there is no graph of maximum degree $k$ or less labelled with only members of $S$ which distinguishes $P$ from $P_{AF}$. From here on we will assume that all unique labels are drawn from $n+\kappa]$ where $\kappa$ is a sufficiently large constant and $n$ is the number of nodes in the graph. We obtain the following result, discussion and sketching the proof of which makes up the remainder of this section, provided $\kappa\geq R(9,8)$.
@@
Let $P=(b,f)$ be a correct and terminating broadcast protocol defined according to definition 4.1, then $P$ is not distinct from Amnesiac Flooding.
@@
The basic argument is to show that any correct and terminating broadcast protocol that meets the criteria is identical to Amnesiac Flooding. Our core technique is to construct a set of network topologies such that any policy distinct from Amnesiac Flooding fails on at least one of them.
For any given protocol we derive a directed graph (separate from the network topology) describing its behaviour and demonstrate via a forbidden subgraph argument that any set of IDs of size $R(x,8)$ must contain a subset $T$ of size at least $x$ that do not respond to each other as leaf nodes, i.e. $P$ is $AF$ up to degree $1$ on $T$. We take $\kappa=R(9,8)$ and show that there must then exist $U\subseteq T$ containing at least $6$ identifiers such that $P$ is AF up to degree $2$ on $U$. By constructing a set of small sub-cubic graphs, we are able to extend this to degree $3$.
These form the base case of a pair of inductive arguments. First, we construct a progression of sub-cubic graphs which enforce that if $P$ is AF on $[m]$ up to degree $3$ it must be AF on $[m+1]$ up to degree $3$. We then construct a family of graphs which have a single node of high-degree, while all other nodes have a maximum degree of $3$ and so must behave as though running Amnesiac Flooding. These graphs permit a second inductive argument showing that this unique high degree node must also behave as if running Amnesiac Flooding. In combination, these two constructions enforce that $P$ behaves like Amnesiac Flooding in all possible cases. The full proof is given in appendix B. ∎
4.2. Relaxing the constraints
Despite the uniqueness established in the previous subsection, we are able to derive four relaxed algorithms distinct from Amnesiac Flooding each obeying only three of the four conditions.
@@ 4.2. Relaxing the constraints
There exist correct and terminating broadcast algorithms distinct from Amnesiac Flooding on infinitely many networks obeying any three of Strict Statelessness, Obliviousness, Determinism and Unit Bandwidth.
The algorithms we obtain all build upon Amnesiac Flooding, and we believe illuminate the role of each of the four conditions in the uniqueness result by showing what they prevent. The algorithms for Obliviousness and Unit Bandwidth are presented together, as they differ only in how control information is encoded.
  • Strict Statelessness: Several relaxations of this already exist, such as Stateless Flooding (the initiator retains information for one round) or even classical non-Amnesiac Flooding (nodes are able to retain 1-bit for one round). We present Neighbourhood-2 Flooding. Nodes know the ID of their neighbours’ neighbours. The protocol behaves distinctly on star graphs, as the hub can determine the entire network topology.
  • Obliviousness and Unit Bandwidth: 1-bit flooding. Nodes are allowed to send a single bit of read-only control information (in the message header or encoded in the number of messages sent) communicating whether the initiator is a leaf vertex. If it is, nodes implement Amnesiac Flooding, otherwise they use a different mechanism called Parrot Flooding (leaves bounce the message back) which always terminates when begun from a non-leaf vertex.
  • Determinism: Random-Flooding. Nodes have access to one bit of randomness per round. Each round every node randomly chooses to implement Amnesiac Flooding or to forward to all neighbours. Random-Flooding is correct with certainty and terminates almost surely in finite time.
Since each of these constitutes only a minor relaxation of the restrictions, we argue that the uniqueness of Amnesiac Flooding is in some sense sharp. We present the protocols fully and demonstrate their correctness and termination for each of these cases independently in appendix C.
@@ 5. Fault Sensitivity
With the uniqueness of Amnesiac Flooding established, a greater understanding of its properties is warranted. In this section, we study the configuration space of Amnesiac Flooding and obtain an exact characterisation of terminating configurations. We then apply this to investigate the algorithm’s fault sensitivity.
5.1. Obtaining a termination dichotomy
@@
A faux-even cycle (FEC) is a graph comprised of either two node disjoint odd cycles connected by a path or two odd cycles sharing only a single node. We denote by $FEC_{x,y,z}$ the $FEC$ with one cycle of length $2x+1$, one of length $2z+1$ and a path containing $y$ edges between them. We emphasize that if $y=0$ the two cycles share a common node and if $y=1$ the two cycles are connected by a single edge.
@@
Let $F=(V,E)$ be $FEC_{x,y,z}$, the even cycle representation of $F$ denoted $F_{2}$ is the graph constructed by splitting the end points of the interconnecting path in two, and duplicating the path to produce an even cycle. Formally, if the two cycles are of the form $a_{0}...a_{2x}a_{0}$ and $c_{0}...c_{2z}c_{0}$, with $a_{0}$ and $c_{0}$ connected by the path $b_{1}...b_{y-1}$, we construct the following large even cycle from four paths:
@@
There then exists a corresponding message configuration over the even cycle representation. Essentially (other than a few technical exceptions), this new configuration is the same as the old configuration but with two copies of each message on the path, one on each corresponding edge of the even cycle representation. Formally,
@@
We first establish that balance, and therefore imbalance, is conserved by Amnesiac flooding and, as the empty configuration is balanced, Amnesiac Flooding cannot terminate from any imbalanced configuration. For Amnesiac Flooding not to terminate it requires that some message travels around the communication graph and returns to the same edge, in the same direction. We show that if a configuration is balanced, the trajectory of any message can spend only a bounded number of consecutive steps on any given cycle or FEC. However, we can also show that any message’s trajectory which crosses the same edge twice in the same direction, must have spent a large number of consecutive steps on some cycle or FEC, and therefore could not have begun in a balanced configuration. Thus, Amnesiac Flooding started from any balanced configuration must terminate. The full proof is given in appendix A. ∎
5.2. Applying the termination dichotomy
In this section, we apply the dichotomy to obtain a number of results.
5.2.1. Extended Dichotomy
While Theorem 5.1 provides a full dichotomy over the configuration space of Amnesiac Flooding and is much easier to reason about than previous results, the definition is somewhat unwieldy. In this subsection, we demonstrate the effectiveness of the dichotomy and unify it with the existing results (Hussak and Trehan, 2023; Turau, 2021a; Hussak and Trehan, 2020b). It has previously been observed that running Amnesiac Flooding backwards obtains another instance of multi-cast Amnesiac Flooding (Hussak and Trehan, 2020b). Formally,
Definition 5.0.
Let $G=(V,E)$ be a graph and $S$ be a configuration of messages on $G$. Then $\bar{S}=\{(v,u)|(u,v)\in S\}$.
Lemma 5.8.
Let $G=(V,E)$ be a graph, $S$ a configuration of messages on $G$ and $T=\{u\in V|\forall v\in N(v):(v,u)\in S\}$ the set of sink vertices. Then $A_{T,G}\overline{A_{\emptyset,G}(\bar{S})}=S$
Which gives the following immediately via induction,
Lemma 5.9.
Let $G=(V,E)$ be a graph and $S$ a configuration of messages on $G$. Then for any $k\in\mathbb{N}$ there exists a sequence $I_{1},...,I_{k}\subseteq V$ such that $A_{I_{1},G}...A_{I_{k},G}\overline{A^{k}_{\emptyset,G}(\bar{S})}=S$.
Intuitively, this means that given any configuration $S$ of Amnesiac Flooding, we can run it backwards through time to some earlier configuration $S^{\prime}$. Further we obtain a sequence of vertex sets $I_{1},...,I_{k}$ that were sinks in the time-reversed process and therefore sources in the forwards process. We can therefore reconstruct $S$ beginning from $S^{\prime}$ via some sequence of fresh multi-casts from $I_{1},...,I_{k}$. We will use this fact to obtain all configurations from which Amnesiac Flooding terminates (i.e. balanced configurations) from the empty configuration. The following lemma is immediate from the definition of balance (definition 5.6), as reversing the direction of all messages in a configuration does not affect its balance.
Lemma 5.10.
$\bar{S}$ is balanced on $G$ if and only if $S$ is balanced on $G$
Putting it all together, we obtain the following extension of the dichotomy result, as well as the complement to Theorem 2.1.
Theorem 5.11.
Let $G=(V,E)$ be a graph and $S$ a configuration of $G$, the following are all equivalent:
  1. (1)
    $\exists k\in\mathbb{N}:A^{k}_{\emptyset,G}(S)=\emptyset$
  2. (2)
    $\exists k\in\mathbb{N},I_{1},...,I_{k}\subseteq V:A_{I_{k},G}...A_{I_{1},G}(\emptyset)=S$
  3. (3)
    $S$ is balanced on $G$
Proof.
The equivalence of (1) and (3) follow immediately from Theorem 5.1. Further, we have that (2) implies (1) from Theorem 2.1. Now assume $S$ is balanced, then by lemma 5.10, so is $\bar{S}$. Thus by Theorem 5.1 there exists a finite $k$ such that after $k$ rounds Amnesiac flooding started from $\bar{S}$ must terminate, i.e. $A^{k}_{\emptyset,G}(\bar{S})=\emptyset$ . Therefore, by lemma 5.9 we have a sequence $I_{1},...,I_{k}\subseteq V$ such that $S=A_{I_{1},G}...A_{I_{k},G}\overline{A^{k}_{\emptyset,G}(\bar{S})}=A_{I_{1},G}...A_{I_{k},G}(\emptyset)$. Thus, we have (3) implies (2) and the result follows. ∎
5.2.2. Fault Sensitivity
@@
Note that we refer to the Byzantine failures as weak, since they are transient and only interfere with the forwarding of the message, not its content. It is obvious to see that in a stateless setting there is no way to deal with a Byzantine fault that changes the message as there is no method to verify which message is authentic. Intuitively, in our setting, Weak-Byzantine agents may choose to send messages to an arbitrary set of neighbours in each round and they are all controlled by a single coordinated adversary. We say that a Weak-Byzantine adversary with control of a given set of nodes can force some behaviour if there exists any weak byzantine failure on that set of nodes producing the forced behaviour. We can now express our fault sensitivity results, the proofs of which we defer to appendix D, and begin with an extreme case of single message dropping.
Theorem 5.12 (Single Message Failure).
If single-source Amnesiac Flooding experiences a single message drop failure for the message $(u,v)$ then it fails to terminate if either:
  • uvuv is not a bridge
  • uvuv lies on a path between odd cycles
Moreover, it fails to broadcast if and only if this is the first message sent along $uv$, uvuv is a bridge, and the side of the cut containing $u$ does not contain an odd cycle.
Thus, Amnesiac Flooding is extremely fault-sensitive with respect to message dropping.
@@
Theorem 5.13 (Uni-directional link failure).
@@
Theorem 5.14 (Byzantine Failure).
@@
  • Non-termination if and only if at least one member of $J$ lies on either a cycle or a path between odd-cycles.
@@ 6. Conclusions and Future Work
We also obtain an understanding of the structural properties of Amnesiac Flooding. In particular, we study its sensitivity to single message drops, uni-directional link failures, and weak byzantine collusion, showing it can easily become non-terminating or non-broadcasting under such conditions. This is perhaps surprising, as statelessness is frequently associated with fault tolerance, such as in the self stabilizing setting. A reasonable interpretation of Theorem 5.1, however, is that Amnesiac Flooding, while locally stateless, depends heavily on a distributed “meta-state” contained in the configuration of sent messages. This suggests it is unlikely that any minor modification of Amnesiac Flooding will resolve its fragility without depending on an entirely different mechanism for termination. In support of this, we note that of the four alternatives presented in the proof of the Existence of Relaxed Protocols Theorem, only Random-Flooding is meaningfully more robust (and will in fact terminate from any configuration in finite time almost surely). Nevertheless, we contend that further exploration of stateless algorithms such as Amnesiac Flooding, their properties and related models are important for both theory and practice of distributed networks.
@@
Appendix A Proof of the Dichotomy (Theorem 5.1)
In this appendix we present all remaining definitions necessary for and prove theorem Theorem 5.1. First, we give the formal definition of the even cycle representation of a message configuration.
With that established, we begin by showing that (im)balance is preserved by the operation of Amnesiac Flooding,
@@
Let $G=(V,E)$ with $H\subseteq G$ an odd cycle of length $2k+1$. If $m$ is a message of $S$ on $H$ with a message path of length $2k+2$ restricted only to $H$, then $S$ is imbalanced on $G$.
@@
For the sake of contradiction assume that $S$ is balanced on $G$ and $m\in S$ has a message path from $S$ of length $2k+2$ restricted only to $H$. Without loss of generality, assume that $m$ is travelling clockwise. On a cycle a message can collide with exactly one other message, upon doing which all of its message paths must terminate. If there is a message $m^{\prime}$ on $H$ travelling anti-clockwise in $S$, then either it collides with $m$ within $2k$ steps or it collides with another message travelling clockwise first. Since $S$ is balanced on $G$ there must be an equal number of messages travelling clockwise and anti-clockwise on $H$, and a collision removes one of each. Therefore, unless new messages are added to $H$ in subsequent steps, $m$ must collide with a message $m^{\prime}$ travelling anti-clockwise before it takes $2k+1$ steps. Thus, new messages must be added to $H$ that collide with the message $m$ would otherwise collide with. The introduction of a new message to $H$ must produce one message travelling clockwise and one travelling anti-clockwise, $m_{c}$ and $m_{a}$ respectively. Thus, $m_{c}$ must collide with $m^{\prime}$ before $m$ does and so must be introduced between $m$ and $m^{\prime}$. However, this means that $m$ will collide with $m_{a}$ before it would have reached $m^{\prime}$ giving our contradiction. ∎
@@
    $W$ cannot return to the node it just came from, i.e. no sequence of the form uvuuvu.
@@
    $W$ cannot take $2x+2$ consecutive steps around an odd cycle of length $2x+1$ (Lemma A.5).
@@
If $W$ contains a factor uuu...u then that factor contains an odd cycle as a subfactor.
@@
Assume there are at least two for contradiction. We take the first two cycles of odd length in $W$, they are either sequential (uuvvu...u...v...v), interleaved (uvuvu...v...u...v) or repetitive (uuuu...u...u). In the sequential case this is a fully traversed $FEC$ and so in violation of Condition $4$. In the interleaved case, we have three subsequences $w_{1},w_{2},w_{3}$ of length $a,b,c$ respectively such that $a+b$ is odd and $b+c$ is odd (see Figure 2). Thus, $a+c$ is even. Since $w_{1}$ and $w_{3}$ induce an even cycle, either $a=c$ or $W$ is in violation of Condition $3$. Therefore, we have that $W$ travels exactly half way around an even cycle along $w_{3}$. Thus there must exist a message path $W^{\prime}$ that contains the factor $w_{1}w_{2}w_{1}$ which is more than $a+b$ steps around an odd cycle of length $a+b$ and so is in violation of Condition $2$. In the repetitive case, this is multiple traversals of an odd cycle in full and so also in violation of Condition $2$. Thus, there exists at most one odd cycle in $W$. ∎
@@
Claim A.14.
$C$ is an odd cycle
Proof:
If $C$ is of even length (say $2k$), then there must be an excursion from $C$ otherwise Condition $3$ would be violated. However, there can be at most one such excursion as it must contain a consecutive odd cycle and so the two subsequences on either side must be factors of $W$. Therefore, since some subsequence of $W$ traverses $C$ fully with one additional step, one of the two factors must take $k+1$ steps around $C$. This violates Condition $3$ and so $C$ must be an odd cycle. $\square$
Claim A.15.
$W$ consists of two odd cycles $C$ and $\hat{C}$ connected by a path.
Proof:
If $C$ is an odd cycle there must be an excursion from it or Condition $2$ would be violated. This excursion must contain exactly one consecutive odd cycle which we denote by $\hat{C}$. If $\hat{C}$ does not share its starting node with $C$, $W$ either forms a path between $C$ and $\hat{C}$ or some chain of cycles. We can use Condition $5$ to eliminate all even cycles of this chain, then if there are any odd cycles in the chain they correspond to a fully traversed $FEC$ when paired with $\hat{C}$ and so violate Condition $4$. Thus, we have a $W$ that takes a simple path from $C$ to $\hat{C}$ and back, although possibly intersecting $C$ along the way. $\square$
Claim A.16.
Neither $C$ nor $\hat{C}$ intersect with the path between them.
Proof:
Assume that the path to $\hat{C}$ does intersect $C$, since we are taking the same path in both directions any node shared between the path and $C$ appear in $W$ three times. This violates Lemma A.12 and so $C$ must be disjoint from the path to $\hat{C}$. Similarly $\hat{C}$ must be disjoint from the path otherwise it would violate the same lemma. If $C$ is disjoint from $\hat{C}$ the pair would form a fully traversed $FEC$, thereby violating Condition $4$. Thus, $C$ and $\hat{C}$ must intersect. $\square$
Claim A.17.
@@
Proof:
Let $W=uv..wx_{1},...,x_{k}w..uv$ where C=uwuvC=u...w...uv and the excursion to $\hat{C}$ is given by $wx_{1}...x_{k}w$. Now assume that $\hat{C}$ contains a node from $C$ which occurs strictly before $w$ in $W$. This node is on a consecutive odd cycle and appears twice. There must exist a latest such node in the ordering of $C$, we denote it $y$. Since $y$ is on a consecutive odd cycle and appears twice in $W$ it must be the start and end point of $\hat{C}$. However, then either $y=w$ which is impossible or $y$ appears three times, violating Lemma A.12. The same argument holds taking the earliest node shared by $C$ and $\hat{C}$ strictly after $w$. Thus, the only node that can be shared by $C$ and $\hat{C}$ is $W$, implying that $W$ forms two odd cycles sharing a single node. However, this is an $FEC$ which is fully traversed and so violates Condition $4$.
Thus, $W$ cannot exist and so there can be no recurrent message in $S$. $\square$ ∎ We can now immediately obtain a proof of corollary 5.2.